Calculate the force applied body on a whose mass 4 is 25ky. The force applied changed the velocity the body from 20m/s to 42 m/s in11 sees


1
Expert's answer
2021-07-05T08:43:04-0400
"F=m\frac{v_f-v_i}{t}\\F=25\frac{42-20}{11}=50\ N"